OpenStudy (anonymous):

If x = K(s, m(s), p(s, t)), find expressions for ∂x/∂s and ∂x/∂t. I got a different solution to another friend of mine. Can somebody see if I am correct or not? My solution: ∂x/∂s = (∂x/∂K)(dK/∂s)+(∂x/∂K)(∂K/∂m)(∂m/∂s)+(∂x/∂K)(∂k/∂p)(∂p/∂s) ∂x/∂t = (∂x/∂K)(∂K/∂p)(∂t/∂p)

OpenStudy (anonymous):

$$\frac{\partial x}{\partial s}=\frac{\partial K}{\partial s}+\frac{\partial K}{\partial m}\cdot\frac{dm}{ds}+\frac{\partial K}{\partial p}\cdot\frac{\partial p}{\partial s}$$this is identical to yours since:$$\frac{dm}{ds}=\frac{\partial m}{\partial s}$$and$$\frac{\partial x}{\partial K}=1$$

OpenStudy (anonymous):

similarly we have$$\frac{\partial x}{\partial t}=\frac{\partial K}{\partial p}\cdot\frac{\partial p}{\partial t}$$

OpenStudy (anonymous):

Thanks so much for this - really appreciate it! Can I ask how you know that ∂x/∂K = 1?

OpenStudy (anonymous):

you said \(x=K\)

OpenStudy (anonymous):

Nice. Hadn't really taken a step back and thought about that relationship fully. Appreciate it a lot man, thanks again!

OpenStudy (anonymous):

np glad I could help